OpenStudy (anonymous):

How would I do this? What is an equation for the line that passes through points (-1, -4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (phi):

do you know how to find the slope between two points?

OpenStudy (anonymous):

Yes, I know how to do that. So, it would be like this. m = y2 - y1 over x2 - x1 m = 4 - (-4) over 1 - (-1) Am I doing it correctly?

OpenStudy (anonymous):

This would make 8 over 2. Therefore, forming 4/1 which turns into 4. @phi: Did I do it right?

OpenStudy (anonymous):

However, they want it in slope-intercept form...

OpenStudy (phi):

yes you found the slope. slope intercept form is y= mx + b where m is the slope and b is the y-intercept (where the line crosses the y-axis) with m=4 you have y= 4x+b you have to find b to finish the problem one way is replace x and y with numbers that are on the line for example (1, 4) is on the line. (1,4) means x=1 and y=4 replace x and y in your equation 4= 4*1+b can you find b?

OpenStudy (anonymous):

Okay, hold on please.

OpenStudy (anonymous):

Why do you have the * symbol after the 4? Would we just put the 1 in parenthesis after it instead?

OpenStudy (anonymous):

B is the y-intercept so that would make it -4 I believe.

OpenStudy (phi):

* means multiply in lots of computer programming languages you could write 4*1 as 4(1). or with a dot \(4\cdot 1\)

OpenStudy (phi):

for example (1, 4) is on the line. (1,4) means x=1 and y=4 replace x and y in your equation 4= 4*1+b or 4= 4+b b=4 would make this 4= 4+4 or 4=8 which is not correct to solve for b, add -4 to both sides 4-4= 4+b-4 and simplify

OpenStudy (phi):

b= -4 is equally bad

OpenStudy (anonymous):

I'm here just to let you know. I had to do something real quick. I'm going to read what you posted now.

OpenStudy (anonymous):

Okay, I understand. They just want me to state the equation for the line in slope-intercept form so it would be this: 4 = 4(1) + -4 Correct?

OpenStudy (phi):

you are missing an idea ? you are asked to find the equation of a line in slope intercept form y= mx +b so the answer will be in that form. notice that 4 = 4(1) + -4 does not have an x or y, so that can't be right. m is the slope. you found the slope. it is m=4 so you know some of the equation y= 4x +b now you need to find b and you will be done.

OpenStudy (anonymous):

Okay, so we just leave the y and x in its variable form like you did, and make sure that we state the values of m and b?

OpenStudy (phi):

yes, now you must find b you use algebra. to find b, use a point you know is on the line. (1,4) is on the line (they told you this) so 4= 4(1) +b must be true 4=4+b what b makes this an equality?

OpenStudy (anonymous):

0?

OpenStudy (phi):

yes. b=0 if you use algebra 4+b= 4 you would write down -4 on both sides of the = 4+b-4= 4 - 4 on the right side 4-4 is 0 4+b-4= 0 on the left, you can change the order when adding: b+4-4 and b+ (4-4) is b+0 or just b we get b=0 so finally, we have your equation y= mx+b with m=4 and b=0 y= 4x +0 but nobody writes + 0, so the answer is y= 4x

OpenStudy (anonymous):

Okay, thanks for your help! :)